# Reinforced Generation of Combinatorial Structures: Hardness of Approximation

### v5

@@ 3.1 Developing a Search Framework for Soundness and Completeness Guarantees
This template requires $k$ separate gadgets $\{I_{i}:0\leq i<k\}$, that correspond to the predicates $\{P_{i}^{=}:0\leq i<k\}$ of $3\textup{LIN}({k})$. An important feature of $\textup{MAX-}{k}\textup{-CUT}$ is its ${\mathbb{S}}_{k}$-symmetry – it is invariant under permuting the alphabet $\Psi=\mathbb{Z}_{k}$. Unfortunately this feature is not shared by $3\textup{LIN}({k})$, and as a result, our reductions require a way to break this symmetry of $\textup{MAX-}{k}\textup{-CUT}$. We achieve this by adding $k$ new global variables to the standard systematization of gadgets [50].
@@
AlphaEvolve found a good $I_{0}$ gadget on $9$ variables (Figure 2(a)) quite quickly, mainly because the symmetries of $P_{0}^{=}$ work well with $\textup{MAX-}{k}\textup{-CUT}$ when $k=3$. In particular, this gadget did not require the aforementioned global variables. Building on ideas from the TSSW framework [50] we can show that this is the optimal $I_{0}$ gadget in terms of the final inapproximability ratio.
In contrast, finding good $I_{1}$ and $I_{2}$ gadgets required significantly more effort and some new ideas. Unlike $I_{0}$, we found that increasing the number of auxiliary variables consistently led to better performing gadgets, with the only bottleneck being the runtime of computing the performance of a candidate gadget (we comment more on this point in Section 3.2).
The $I_{1}$ gadget we use in the proof of Theorem 3.1 (Figure 2(b)) requires $14$ total variables. We have some limited evidence that our $I_{1}$ gadget is “locally” optimal in the sense that it gives the optimal inapproximability result among all gadgets on $14$ variables with the same witness as $I_{1}$ for all but one satisfying assignment $\mathbf{x}$ of $P_{1}^{=}$ (see the completeness case of Definition C.1 for what is meant by a “witness”) [^6] Such a gadget giving the optimal inapproximability result can be quantified as a mixed integer program, which we solved computationally to determine that our gadget is the optimal one. See Section 3.3 for more details..
@@
The main difference in our process of finding gadgets for $\textup{MAX-}{4}\textup{-CUT}$ was that we had to search over gadgets with more variables (as many as $19$), so we required an extremely well-performing implementation of the verifier (again, we elaborate in Section 3.2). Even with this optimized verifier, evaluations were quite slow, requiring on the order of one second to evaluate a single gadget on $19$ variables. As a result, it took a lot longer for AlphaEvolve to find a search algorithm that, given query access to this optimized verifier, could find any nontrivial gadget. The final search algorithm AlphaEvolve produced had the distinctive property of searching over weighted gadgets with real-valued weights, resulting in gadgets with a wide variety of weights between $1$ and $1429$ after appropriate scaling and rounding.
@@ 3.2 Faster Verification via AlphaEvolve to Explore Larger Gadgets
The main challenge with finding large gadgets is that the cost of scoring a gadget scales exponentially in the number of variables; computing the completeness and soundness parameters described in Definition C.1 essentially amount to solving an instance of $\textup{MAX-}{k}\textup{-CUT}$, which requires exponential time. Even for $k=3$, AlphaEvolve slows down significantly when searching for gadgets of size as few as $11$ with a brute force $\textup{MAX-}{3}\textup{-CUT}$ solver.
This problem does not have an off-the-shelf solution in the form of existing fast solvers; it is unlikely that existing solvers [17, 41] can be repurposed to solve $\textup{MAX-}{k}\textup{-CUT}$. To solve this issue, we used AlphaEvolve itself to speed up a naive brute force implementation of $\textup{MAX-}{k}\textup{-CUT}$, scoring a candidate implementation by runtime and correctness. In order to calculate the runtime, we created a synthetic dataset of a wide variety of $\textup{MAX-}{k}\textup{-CUT}$ instances, drawn from $20$ random models with varying amounts of planted structure. We then tasked AlphaEvolve to maximize the number of variables $m$ for which the solver requires at most one second on average to solve instances of size $m$ from our dataset.
The biggest challenge was ensuring that AlphaEvolve does not cheat and find an incorrect solver that is much faster. As mentioned before, we achieved correctness by (1) checking that the solver is correct on our synthetic dataset, and (2) using a separate judge LLM to certify that a candidate solver is corr ect. Each of these techniques was individually too lenient to avoid incorrect solvers, but we found by human inspection that they were enough in combination.
We note that once $m$ is large enough, it is not possible to label our dataset with the “ground truth” scores using a brute-force implementation (which is guaranteed to be correct). Instead, we inductively rely on the correctness of previous solvers produced by AlphaEvolve (which have already passed the above correctness checks) that are fast enough to provide labels for large $m$.
We did this separately for the $k=3$ and $k=4$ cases, obtaining solvers that are optimized to each particular problem. We found that systems-level improvements were most important to the $k=3$ case, with the final solver offloading the main $O(3^{m})$ time computation to a highly optimized tensor contraction operation in numpy, and only performing slow python-based computation for $O(3^{m/3})$ time. This solver resulted in a $10,000\times$ speedup for instances of size $m=14$.
For $k=4$, the final solver used a more sequential branch-and-bound strategy along with some systems-level improvements. At $m=19$, this provided a $10,000\times$ speedup against even a numba-accelerated [39] brute force solver.
@@ 3.3 Comparison to other computational techniques
The most straightforward comparison is the TSSW framework [50], which casts the task of finding the optimal gadget as a linear program (LP). The main difficulty in doing this is the presence of existential quantifiers in computing the completeness of the gadget. In order to eliminate these quantifiers, [50] canonicalize the auxiliary variables in the gadget. As a result, the size of the LP encoding the optimal gadget is doubly exponential in the number of satisfying assignments of the source predicate; this is $3^{3^{6}}$ for a reduction from $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$, which is computationally infeasible. Sometimes (as is the case for our $I_{0}$ gadget for $\textup{MAX-}{3}\textup{-CUT}$) it is possible to argue that not all auxiliary variables are required, leading to a more tractable LP, but it is not clear that this is possible outside of very special cases.
@@ 4 NP-Hardness of Approximating Metric TSP
The Traveling Salesman Problem (TSP) [53] is one of the most studied problems in combinatorial optimization. In this paper we consider the metric version of the problem, where the objective is to find a minimum weight Hamiltonian cycle in a weighted, complete, undirected graph where the weights satisfy the triangle inequality. The best approximation algorithm achieves a factor of $1.5-10^{-34}$ [20]. A series of results [45, 16, 7, 44, 37, 33, 11] led up to the SOTA NP-hardness of approximation to a factor of $117/116$. We use AlphaEvolve to improve this result to $111/110$.
@@
A sparse CSP instance: We follow the framework of [37, 33, 11] and start with an NP-hard CSP we call $\text{Hybrid-}3\textup{LIN}({2})$, which we reduce to metric TSP to obtain Theorem 4.1. $\text{Hybrid-}3\textup{LIN}({2})$ is a particularly structured weighted CSP where every variable appears in exactly three equations, where the equations are either $3\textup{LIN}({2})$ constraints of the form $x\oplus y\oplus z=b$, or $2$-ary linear constraints of the form $x\oplus y=b$, for some $b\in\{0,1\}$. One can show that given a $\text{Hybrid-}3\textup{LIN}({2})$ instance, it is NP-hard to decide whether (1) most equations can be satisfied, or (2) one cannot satisfy a significant fraction of the equations. (This is the same source problem in [11], and a formal statement about the hardness of approximation is provided in Theorem D.1.)
A spanning tour in a graph is defined as a closed walk that visits each vertex at least once. It is convenient to work with an equivalent formulation of metric TSP [33] (in terms of inapproximability), which we denote by MWST. The goal there is to find the minimum weight spanning tour in a weighted, undirected graph $G=(V,E,w)$. [11] first reduces from an instance of $3\textup{LIN}({2})$ to an instance of $\text{Hybrid-}3\textup{LIN}({2})$ via an expander graph based CSP-to-CSP reduction. Then they reduce from $\text{Hybrid-}3\textup{LIN}({2})$ to MWST.
@@
Reducing the sparse CSP to MWST: In the gadget reduction from $\text{Hybrid-}3\textup{LIN}({2})$ to MWST each variable is replaced by a vertex, and each equation is replaced by a corresponding gadget. An equation gadget (more concretely described below) gives a way to encode an assignment of True / False to the variables in a $3\textup{LIN}({2})$ equation into a connectivity property for the corresponding vertices, in the sense that if the equation is satisfied, one can connect the involved vertices to the rest of the graph with low cost; on the other hand, if the equation is unsatisfied, then any spanning tour must dedicate a large total weight of edges in order to connect the involved vertices to the rest of the graph.
Using AlphaEvolve, we find a new equation gadget (Figure 4) achieving better performance than the one appearing in [11] (Figure 8(a)) [^8] The $3\textup{LIN}({2})$ predicate encoded in the equation gadgets in [11, 33, 37] all have the form $x\oplus y\oplus z=0$, as opposed to $x\oplus y\oplus z=1$ in Figure 4.. In order to quantify this improvement, we now describe equation gadgets more concretely: each $3\textup{LIN}({2})$ predicate (of the form $x\oplus y\oplus z=1$) is assigned an equation gadget, where the green vertices $\{1,2,3\}$ in Figure 4, also called the contact vertices, correspond to the variables $\{x,y,z\}$. The red vertex $\{4\}$ is called the central vertex, and is shared across all appearances of the equation gadget across all $3\textup{LIN}({2})$ equations in the instance. The rest of the vertices in the equation gadget (shown in blue in Figure 2) are used to ensure that there is a gap between the weight of a spanning tour due to a satisfied $3\textup{LIN}({2})$ equation versus an unsatisfied one. The black edges (referred to as “unforced edges”) are optional in any tour, while any tour is forced to take each of the red edges (referred to as “forced edges”) at least once. We note that a standard trick (e.g., [33]) can be used to implement the constraint of forcing certain edges to be taken within MWST. The green dashed edges are called special edges, and are purely used for analysis purposes, and do not appear in the actual MWST instance.
@@
Figure 4: Equation gadget for MWST found by AlphaEvolve, when reducing from a $3\textup{LIN}({2})$ equation of the form $x\oplus y\oplus z=1$. Vertices $\{1,2,3\}$ represent variables in the $3\textup{LIN}({2})$ equation. The red edges represent the forced edges. The dashed green edges represent the special edges. All red and black edges have weight one.
@@
The optimization problem for the equation gadget search can be cast as a mixed integer program (MIP). Because of its simplicity, it is perceivable that the gadget in Figure 4 can be found by directly solving the MIP assuming the number of auxiliary vertices are known in advance. However, because of the large number of constraints involved in the equation gadget corresponding to $x\oplus y\oplus z=0$ (in Figure 8(b)), i.e., around $11!$, using traditional solvers would face the same computational bottlenecks described (in Section 3.3) for the $\textup{MAX-}{k}\textup{-CUT}$ problem.
@@ 5 Discussion on AI-assisted Mathematics and Complexity Theory
We begin this section with a discussion of the norms by which to understand how AI can assist research in the mathematical sciences. We then discuss the role of AlphaEvolve in achieving the results in this paper, highlighting how our work differs from prior work in computer-assisted theorem proving.
In studying the role of AI in assisting mathematical discovery, we must consider at least four scenarios:
@@
    A human invokes a language model to summarize the state of prior art, to chart a research plan towards new theorems, or to directly generate fragments of (or entire) proofs.
@@
    A human uses AI-derived tools such as AlphaEvolve to generate better proof elements (gadgets, graphs).
@@
    A human uses custom code independent of AI to discover better proof elements.
@@
    A human discovers the same or better proof elements by hand.
@@
The idea that LLMs can usefully generate summaries of prior literature in a field has been mooted about for some time [52, 54, 29, 22], and indeed our experience confirmed this on a number of prompts. The results (modulo occasional hallucination) are a good starting point for deeper exploration and understanding. Interestingly, across many examples we could rapidly generate an overview of the art in an unfamiliar field; we believe this capability will increasingly be used by scientists working across fields so that — for instance — an algebraist can come up to speed with Ramsey Theory. We believe that in time, this will lead to more fluent cross-pollination across disciplines. We have not yet been able to prompt an LLM into providing a usable research plan to obtain new results (such as the ones we report), and this is the focus of deeper efforts e.g., Google’s Co-Scientist [24] program.
@@
There have been prior efforts to generate proofs for open mathematical statements via prompting an LLM directly [46, 10, 43, 13, 31]. However, this approach has seen limited success. In some cases an LLM was indeed able to generate the complete and correct proof of a previously unproven statement [10] (although it is possible that a persistent human could perhaps have derived this), but in many cases [46, 43, 13, 31] the LLM could only generate a proof sketch, which ultimately had to be filled in by humans. In general, these proofs/sketches currently require a human to verify correctness. Indeed, our natural attempts to prompt a standard LLM into directly generating the kinds of combinatorial structures in this paper met with failure  [^9] For example, GPT5 and Gemini 2.5 Pro could not generate the $16/17$-NP-hardness gadget for MAX-CUT, even though the gadget is explicitly mentioned in [50, Figure 4.1] — the prompt being Can you generate the gadget that gives the construction for $16/17$ approximation of MAX-CUT under NP hardness?. While this may become feasible as LLM reasoning capabilities improve [38], a formal comparison is beyond the current scope. We emphasize that our AlphaEvolve constructions always come with a certificate of correctness that is formally verified via standard computational approaches. Once the verification code is sound, there is no further need for human scrutiny.
Next we discuss whether a human could have discovered the combinatorial structures in our proofs by hand. It is conceivable that a human expert or a highly customized computational solver could eventually find these solutions. However, their discovery is likely beyond the reach of simple ”pencil-and-paper” methods, and standard SMT/MIP solvers failed to produce them.
@@
In both average case hardness for random graphs, and NP-hardness for approximating $\textup{MAX-}{k}\textup{-CUT}$, computer assisted methods [50, 26, 36] have been used previously. [36] used computational methods to generate $d$-regular Ramanujan graphs (for $d\in\{3,4\}$) with large cut values (or independent sets). While they do not specify their computational approach, we could replicate their results by random sampling of $d$-regular graphs, and testing for the properties by brute force. For reasons mentioned in Section 2[36] could demonstrate their lower bounds for $n\leq 12$, whereas the graphs we find go up to $n=163$.
@@
For the NP-hardness gadget reduction in Section 3, both the soundness and the completeness constraints can be translated into a linear program via skolemization [50]. However, the size of such a linear program is doubly exponential in the number of vertices in the constraint graph. As a result, even for $\textup{MAX-}{3}\textup{-CUT}$, running a linear program (LP) with the canonical number of variables [50] $\approx 3^{3^{6}}$ becomes infeasible with the standard LP solvers we know of. One can directly attack the non-convex program using MIP/SMT solvers, but with the number of constraints being exponential in the number of variables $n\geq 14$, they too did not seem to scale. (See Section 3.3 for a detailed discussion.) In the case of hardness of approximating the TSP, the soundness and completeness constraints can also be cast as MIP constraints. Even with a modest number of vertices in the equation gadget (e.g., twelve in Figure 8(b)), the number of constraints becomes prohibitively large for standard MIPs to handle (around $11!\approx 3\times 10^{7}$).

### v6

@@ 3.1 Developing a Search Framework for Soundness and Completeness Guarantees
This template requires $k$ separate gadgets $\{\texttt{I}^{\equiv i}_{3,k}:0\leq i<k\}$, that correspond to the predicates $\{\texttt{P}^{\equiv i}_{3,k}:0\leq i<k\}$ of $3\textup{LIN}({k})$. An important feature of $\textup{MAX-}{k}\textup{-CUT}$ is its ${\mathbb{S}}_{k}$-symmetry – it is invariant under permuting the alphabet $\mathbb{Z}_{k}$. Unfortunately this feature is not shared by $3\textup{LIN}({k})$, and as a result, our reductions require a way to break this symmetry of $\textup{MAX-}{k}\textup{-CUT}$. We achieve this by adding $k$ new global variables to the standard systematization of gadgets [59].
@@
AlphaEvolve found a good $\texttt{I}^{\equiv 0}_{3,3}$ gadget on $9$ variables (Figure 2(a)) quite quickly, mainly because the symmetries of $\texttt{P}^{\equiv 0}_{3,k}$ work well with $\textup{MAX-}{k}\textup{-CUT}$ when $k=3$. In particular, this gadget did not require the aforementioned global variables. Building on ideas from the TSSW framework [59] we can show that this is the optimal $\texttt{I}^{\equiv 0}_{3,3}$ gadget in terms of the final inapproximability ratio.
In contrast, finding good $\texttt{I}^{\equiv 1}_{3,3}$ and $\texttt{I}^{\equiv 2}_{3,3}$ gadgets required significantly more effort and some new ideas. Unlike $\texttt{I}^{\equiv 0}_{3,3}$, we found that increasing the number of auxiliary variables consistently led to better performing gadgets, with the only bottleneck being the runtime of computing the performance of a candidate gadget (we comment more on this point in Section 3.2).
The $\texttt{I}^{\equiv 1}_{3,3}$ gadget we use in the proof of Theorem 3.1 (Figure 2(b)) requires $14$ total variables. We have some limited evidence that our $\texttt{I}^{\equiv 1}_{3,3}$ gadget is “locally” optimal in the sense that it gives the optimal inapproximability result among all gadgets on $14$ variables with the same witness as $\texttt{I}^{\equiv 1}_{3,3}$ for all but one satisfying assignment $\mathbf{x}$ of $\texttt{P}^{\equiv 1}_{3,k}$ (see the completeness case of Definition C.1 for what is meant by a “witness”) [^6] Such a gadget giving the optimal inapproximability result can be quantified as a mixed integer program, which we solved computationally to determine that our gadget is the optimal one. See Section 3.3 for more details..
@@
The main difference in our process of finding gadgets for $\textup{MAX-}{4}\textup{-CUT}$ was that we had to search over gadgets with more variables (as many as $19$), so we required an extremely well-performing implementation of the verifier (again, we elaborate in Section 3.2). Even with this optimized verifier, evaluations were quite slow, requiring on the order of one second to evaluate a single gadget on $19$ variables. As a result, it took a lot longer for AlphaEvolve to find a search algorithm that, given this optimized verifier, could find any nontrivial gadget. The final search algorithm AlphaEvolve produced had the distinctive property of searching over weighted gadgets with real-valued weights, resulting in gadgets with a wide variety of weights between $1$ and $1429$ after appropriate scaling and rounding.
@@ 3.2 Faster Verification via AlphaEvolve to Explore Larger Gadgets
The main challenge with finding large gadgets is that the cost of scoring a gadget scales exponentially in the number of variables; computing the completeness and soundness parameters described in Definition C.1 essentially amount to solving an instance of $\textup{MAX-}{k}\textup{-CUT}$, which requires exponential time. Even for $k=3$, AlphaEvolve slows down significantly when searching for gadgets of size as few as $11$ with a brute force $\textup{MAX-}{3}\textup{-CUT}$ verifier.
This problem does not have an off-the-shelf solution in the form of existing fast verifiers; it is unlikely that existing SMT/MIP solvers [21, 47] can be repurposed to solve $\textup{MAX-}{k}\textup{-CUT}$. To solve this issue, we used AlphaEvolve itself to speed up a naive brute force implementation of $\textup{MAX-}{k}\textup{-CUT}$, scoring a candidate implementation by runtime and correctness. In order to calculate the runtime, we created a synthetic dataset of a wide variety of $\textup{MAX-}{k}\textup{-CUT}$ instances, drawn from $20$ random models with varying amounts of planted structure. We then tasked AlphaEvolve to maximize the number of variables $m$ for which the verifier requires at most one second on average to solve instances of size $m$ from our dataset.
The biggest challenge was ensuring that AlphaEvolve does not cheat and find an incorrect verifier that is much faster. As mentioned before, we achieved correctness by (1) checking that the verifier is correct on our synthetic dataset, and (2) using a separate judge LLM to certify that a candidate verifier is correct. Each of these techniques was individually too lenient to avoid incorrect verifiers, but we found by human inspection that they were enough in combination.
We note that once $m$ is large enough, it is not possible to label our dataset with the “ground truth” scores using a brute-force implementation (which is guaranteed to be correct). Instead, we inductively rely on the correctness of previous verifiers produced by AlphaEvolve (which have already passed the above correctness checks) that are fast enough to provide labels for large $m$.
We did this separately for the $k=3$ and $k=4$ cases, obtaining verifiers that are optimized to each particular problem. We found that systems-level improvements were most important to the $k=3$ case, with the final verifier offloading the main $O(3^{m})$ time computation to a highly optimized tensor contraction operation in numpy, and only performing slow python-based computation for $O(3^{m/3})$ time. This verifier resulted in a $10,000\times$ speedup for instances of size $m=14$.
For $k=4$, the final verifier used a more sequential branch-and-bound strategy along with some systems-level improvements. At $m=19$, this provided a $10,000\times$ speedup against even a numba-accelerated [45] brute force verifier.
@@ 3.3 Comparison to other computational techniques
The most straightforward comparison is the TSSW framework [59], which casts the task of finding the optimal gadget as a linear program (LP). The main difficulty in doing this is the presence of existential quantifiers in computing the completeness of the gadget. In order to eliminate these quantifiers, [59] canonicalize the auxiliary variables in the gadget. As a result, the size of the LP encoding the optimal gadget is doubly exponential in the number of satisfying assignments of the source predicate; this is $3^{3^{6}}$ for a reduction from $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$, which is computationally infeasible. Sometimes (as is the case for our $\texttt{I}^{\equiv 0}_{3,3}$ gadget for $\textup{MAX-}{3}\textup{-CUT}$) it is possible to argue that not all auxiliary variables are required, leading to a more tractable LP, but it is not clear that this is possible outside of very special cases.
@@ 4 NP-Hardness of Approximating Metric TSP
The Traveling Salesman Problem (TSP) [64] is one of the most studied problems in combinatorial optimization. While there are many variants of this problem [38, 57, 56, 61], we specifically focus on the metric variant of the problem [38], where the objective is to find a minimum weight Hamiltonian cycle in a weighted, complete, undirected graph with the weights satisfying the triangle inequality. This in turn corresponds to outputting a permutation over the vertices in the graph that has the minimum weight cycle. In this paper we are concerned with the hardness of approximating metric TSP. The best approximation algorithm achieves a factor of $1.5-10^{-34}$ [25]. A series of results [51, 20, 10, 50, 43, 39, 15] led up to the SOTA NP-hardness of approximation to a factor of $117/116$. We use AlphaEvolve to improve this result to $111/110$.
@@
A sparse CSP instance: We follow the framework of [43, 39, 15] and start with an NP-hard CSP we call $\text{Hybrid-}3\textup{LIN}({2})$, which we reduce to metric TSP to obtain Theorem 4.1. $\text{Hybrid-}3\textup{LIN}({2})$ is a particularly structured weighted CSP where every variable appears in exactly three equations, where the equations are either $3\textup{LIN}({2})$ constraints of the form $x\oplus y\oplus z=b$, or binary linear constraints of the form $x\oplus y=b$, for some $b\in\{0,1\}$. One can show that given a $\text{Hybrid-}3\textup{LIN}({2})$ instance, it is NP-hard to decide whether (1) most equations can be satisfied, or (2) one cannot satisfy a significant fraction of the equations. (This is the same source problem in [15], and a formal statement about the hardness of approximation is provided in Theorem D.1.)
A spanning tour in a graph is defined as a closed walk that visits each vertex at least once. It is convenient to work with an equivalent formulation of metric TSP [39] (in terms of inapproximability), which we denote by MWST. The goal there is to find the minimum weight spanning tour in a weighted, undirected graph $G=(V,E,w)$. [15] first reduces from an instance of $3\textup{LIN}({2})$ to an instance of $\text{Hybrid-}3\textup{LIN}({2})$ via an expander graph based CSP-to-CSP reduction. Then they reduce from $\text{Hybrid-}3\textup{LIN}({2})$ to MWST. Intuitively, the equivalence stems from the triangle inequality, which guarantees that any closed walk visiting vertices multiple times can be “short-cut” to form a valid Hamiltonian cycle of equal or lesser cost.
This MWST formulation is particularly advantageous for hardness reductions. By shifting the focus from finding a simple cycle to finding a minimum-cost connected Eulerian subgraph, one avoids the cumbersome task of explicitly constructing the metric closure. This allows the reduction to define weights locally on the sparse graph—typically assigning small weights to edges and large penalties to non-edges—while ensuring that the structural properties of the original hard instance are preserved.
@@
Reducing the sparse CSP to MWST: In the gadget reduction from $\text{Hybrid-}3\textup{LIN}({2})$ to MWST each variable is replaced by a vertex, and each equation is replaced by a corresponding gadget. An equation gadget (more concretely described below) gives a way to encode an assignment of True / False to the variables in a $3\textup{LIN}({2})$ equation (i.e., clauses of the form $\texttt{P}^{\equiv 0}_{3,2}$ or $\texttt{P}^{\equiv 1}_{3,2}$) into a connectivity property for the corresponding vertices, in the sense that if the equation is satisfied, one can connect the involved vertices to the rest of the graph with low cost; on the other hand, if the equation is unsatisfied, then any spanning tour must dedicate a large total weight of edges in order to connect the vertices involved to the rest of the graph.
Using AlphaEvolve, we find a new equation gadget (Figure 4) achieving better performance than the one appearing in [15] (Figure 8(a)) [^8] The $3\textup{LIN}({2})$ equation encoded in the equation gadgets in [15, 39, 43] all have the form $x\oplus y\oplus z=0$, as opposed to $x\oplus y\oplus z=1$ in Figure 4.. In order to quantify this improvement, we now describe equation gadgets more concretely: each $3\textup{LIN}({2})$ equation $\texttt{P}^{\equiv 1}_{3,2}$ (of the form $x\oplus y\oplus z=1$) is assigned an equation gadget, where the green vertices $\{1,2,3\}$ in Figure 4, also called the contact vertices, correspond to the variables $\{x,y,z\}$. The red vertex $\{4\}$ is called the central vertex, and is shared across all appearances of the equation gadget across all $3\textup{LIN}({2})$ equations in the instance. The rest of the vertices in the equation gadget (shown in blue in Figure 2) are used to ensure that there is a gap between the weight of a spanning tour due to a satisfied $3\textup{LIN}({2})$ equation versus an unsatisfied one. The black edges (referred to as “unforced edges”) are optional in any tour, while any tour is forced to take each of the red edges (referred to as “forced edges”) at least once. We note that a standard trick (e.g., [39]) can be used to implement the constraint of forcing certain edges to be taken within MWST. The green dashed edges are called special edges, and are purely used for analysis purposes, and do not appear in the actual MWST instance.
@@
Figure 4: Equation gadget for MWST found by AlphaEvolve, when reducing from a $3\textup{LIN}({2})$ equation $\texttt{P}^{\equiv 1}_{3,2}$ of the form $x\oplus y\oplus z=1$. Vertices $\{1,2,3\}$ represent variables in the $3\textup{LIN}({2})$ equation. The red edges represent the forced edges. The dashed green edges represent the special edges. All red and black edges have weight one.
@@
The optimization problem for the equation gadget search can be cast as a mixed integer program (MIP). Because of its simplicity, it is perceivable that the gadget in Figure 4 can be found by directly solving the MIP assuming the number of auxiliary vertices are known in advance. However, because of the large number of constraints involved in the equation gadget corresponding to $x\oplus y\oplus z=0$ (in Figure 8(b)), i.e., around $11!$, using traditional SMT/MIP solvers would face the same computational bottlenecks described (in Section 3.3) for the $\textup{MAX-}{k}\textup{-CUT}$ problem.
@@ 5 Discussion on AI-assisted Mathematics and Complexity Theory
In studying the role of AI in assisting mathematical discovery, we must consider at least these scenarios:
@@
    We invoke a language model to summarize the state of prior art, to chart a research plan towards new theorems, or to directly generate fragments of (or entire) proofs.
@@
    We use AI-derived tools such as AlphaEvolve to generate better proof elements (gadgets, graphs).
@@
    We use custom code independent of AI to discover better proof elements.
@@
    We discover the same or better proof elements by hand.
  5. 5.
    A combination of (1)-(4).
@@
Literature summaries: The idea that LLMs can usefully generate summaries of prior literature in a field has been mooted about for some time [63, 65, 34, 27], and indeed our experience confirmed this on a number of prompts. The results (modulo occasional hallucination) are a good starting point for deeper exploration and understanding. Interestingly, across many examples we could rapidly generate an overview of the art in an unfamiliar field; we believe this capability will increasingly be used by scientists working across fields so that — for instance — an algebraist can come up to speed with Ramsey Theory. We believe that in time, this will lead to more fluent cross-pollination across disciplines. We have not yet been able to prompt an LLM into providing a usable research plan to obtain new results (such as the ones we report), and this is the focus of deeper efforts e.g., Google’s Co-Scientist [29] program.
Direct Prompting: There have been efforts to generate proofs for open mathematical statements via prompting an LLM directly [52, 14, 49, 17, 36, 9]. This approach has seen mixed success. In some cases an LLM was indeed able to generate the complete and correct proof of a previously unproven statement [14, 9] (although it is possible that a persistent human could perhaps have derived the same), but in many cases [52, 49, 17, 36] the LLM could only generate a proof sketch, which ultimately had to be filled in by humans. At the time of this writing, one significant piece of work in this category is [9]. This report documents GPT-5’s ability to accelerate research in mathematics and theoretical computer science, demonstrating how “scaffolding” — the technique of priming the model with simpler, related warm-up problems — enabled it to derive improved bounds for online algorithms and formally prove previously open conjectures in graph theory. The authors highlight the model’s capacity to go beyond retrieval, successfully constructing novel counterexamples and proof strategies that had previously eluded human experts, such as generating a complex counterexample to the “Follow-the-Leader” algorithm in convex body chasing and proposing the “stability-style analysis” key to solving Erdős Problem #848 [12].
@@
In general, these proofs/sketches currently require a human to verify correctness. Our own natural attempts to prompt a standard LLM into directly generating the kinds of combinatorial structures in this paper met with failure  [^9] For example, GPT 5.2 and Gemini 3.0 Ultra could not generate the $16/17$-NP-hardness gadget for MAX-CUT, even though the gadget is explicitly mentioned in [59, Figure 4.1] — the prompt being Can you generate the gadget that gives the construction for $16/17$ approximation of MAX-CUT under NP hardness?. While this may become feasible as LLM reasoning capabilities improve [44], a formal comparison is beyond the current scope. We emphasize that our constructions always come with a certificate of correctness that is formally verified via standard computational approaches. Once the verification code is sound, there is no further need for human scrutiny.
Demonstrating AlphaEvolve’s breadth: The recent paper [24] is a sweeping tour de force of LLM-guided evolutionary search (AlphaEvolve) in autonomously discovering novel mathematical constructions across analysis, combinatorics, and geometry, surprisingly finding a counterexample to the “Four Guards” logic puzzle by creatively engaging in “prompt injection” against the verifying language model. It notably improves bounds on long-standing problems like the Kakeya needle problem and the Ring Loading Problem, often achieving results competitive with or superior to human-designed benchmarks with minimal problem-specific tuning.
@@
Computational methods in hardness of approximation: In both average case hardness for random graphs, and NP-hardness for approximating $\textup{MAX-}{k}\textup{-CUT}$, computer assisted methods [59, 31, 42] have been used previously. [42] used computational methods to generate $d$-regular Ramanujan graphs (for $d\in\{3,4\}$) with large cut values (or independent sets). While they do not specify their computational approach, we could replicate their results by random sampling of $d$-regular graphs, and testing for the properties by brute force. For reasons mentioned in Section 2[42] could demonstrate their lower bounds for $n\leq 12$, whereas the graphs we find go up to $n=163$.
@@
For the NP-hardness gadget reduction in Section 3, both the soundness and the completeness constraints can be translated into a linear program via skolemization [59]. However, the size of such a linear program is doubly exponential in the number of vertices in the constraint graph. As a result, even for $\textup{MAX-}{3}\textup{-CUT}$, running a linear program (LP) with the canonical number of variables [59] $\approx 3^{3^{6}}$ becomes infeasible with the standard LP solvers we know of. One can directly attack the non-convex program using SMT/MIP solvers, but with the number of constraints being exponential in the number of variables $n\geq 14$, they too did not seem to scale. (See Section 3.3 for a detailed discussion.) In the case of hardness of approximating the TSP, the soundness and completeness constraints can also be cast as MIP constraints. Even with a modest number of vertices in the equation gadget (e.g., twelve in Figure 8(b)), the number of constraints becomes prohibitively large for standard MIPs to handle (around $11!\approx 3\times 10^{7}$).
Gadget design by hand: It is conceivable that a human expert or a highly customized computational verifiers could eventually find these solutions. However, their discovery is likely beyond the reach of simple ”pencil-and-paper” methods, and standard SMT/MIP solvers failed to produce them. Further, humans intuitively cut the search space through insights into symmetries in the constructed objects; in the case of our TSP gadget, it appears that asymmetry was central to the improvement obtained.
AI with significant human effort: Our work and [58] are instances where AI and some significant human effort were deployed in combination. In particular, [58] chronicles the rapid resolution of Erdős Problem #1026 [12] through a hybrid workflow where human ingenuity guided multiple AI systems—including AlphaEvolve for generating optimal constructions and the automated theorem prover Aristotle [2] for formal verification in Lean. By integrating crowd-sourced mathematical insights with AI-driven deep literature search and computational discovery, the collaboration successfully generated the solution in under 48 hours. In our case, we had to refactor (by hand) the proof logic especially for metric TSP, in order for the problem to be amenable to AlphaEvolve.